OpenStudy (anonymous):

-y/8 + 1 = -5 What is the value of y?

OpenStudy (anonymous):

its - y over 8 for any who didn't know

OpenStudy (anonymous):

subtract 1 from both sides then multiply both sides by \(-1\)

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

You can either get a common denominator (8), or you could move the one over and then cross multiply.

OpenStudy (anonymous):

\[\frac{ -y }{ 8 }+\frac{ 1 }{ 1 }=\frac{ -5 }{ 1 }\]

OpenStudy (anonymous):

satellite does it equal 6?

OpenStudy (anonymous):

there is no such mathematical operations as "move" neither is there "cross multiply

OpenStudy (anonymous):

\[\frac{ -y }{ 8 }=\frac{ -6 }{ 1 }\]

OpenStudy (anonymous):

You subtract 1 from each side and then you multiply both sides by 8 and then -y=?

OpenStudy (anonymous):

no \[-\frac{y}{8}+1=-5\] subtract \(1\) get \[-\frac{y}{8}=-6\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then multiply both sides by \(-8\)

OpenStudy (anonymous):

i meant "no, it is not 6"

OpenStudy (anonymous):

and also where did the 1 come from if you got rid of it?

OpenStudy (anonymous):

\[-y=-42\] \[y=42\]

OpenStudy (anonymous):

the 1 was in the question got rid of it by subtracting 1 from both sides of the equation

OpenStudy (anonymous):

is it 48?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thank you very much!